OpenStudy (anonymous):

Find the domain of a function: G(x)=8/8-5x

OpenStudy (mertsj):

all real numbers not equal to 8/5

jimthompson5910 (jim_thompson5910):

You cannot divide by zero, so 8-5x can't be zero If it were, then 8-5x = 0 means that x = 8/5 So when x = 8/5, the denominator 8-5x is zero So this means that we must remove this value from the domain. So the domain is the set of all real numbers except x cannot equal 8/5

OpenStudy (anonymous):

x is a real number and x\[\neq 0\]

jimthompson5910 (jim_thompson5910):

close, more like x is a real number and \[\Large x \neq \frac{8}{5}\]

OpenStudy (anonymous):

oh ok...i got confused with the equal sign with that slash through it...thank you Jim

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

question. so with this i would have the open signs on each end correct ( )?

jimthompson5910 (jim_thompson5910):

are you referring to using interval notation?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

yes, you would use parenthesis and the domain in interval notation is \[\Large \left(-\infty,\frac{8}{5}\right)\cup\left(\frac{8}{5},\infty\right)\]

OpenStudy (anonymous):

thats is what i got, however, i am a little confused on one part..i have to choose the right domain and this is what i have. \[ a. [x \ge8/5], b. {x \neq8}, c. [x \neq8/5], or d. {x \neq0}\]

jimthompson5910 (jim_thompson5910):

it's choice C

OpenStudy (anonymous):

c would be my answer correct?

jimthompson5910 (jim_thompson5910):

yep, you nailed it

jimthompson5910 (jim_thompson5910):

basically x can be any number but 8/5

OpenStudy (anonymous):

thats what i thought..i did have my answer, just confused sometime with those signs { } and [ ].

OpenStudy (anonymous):

thanks bunches again. aloha

jimthompson5910 (jim_thompson5910):

well {} refers to sets

jimthompson5910 (jim_thompson5910):

and [] is used in interval notation

jimthompson5910 (jim_thompson5910):

C is the closest even though the [] should be {}

OpenStudy (anonymous):

oh ok, i will write that for my notes, i am new to this algebra1, never taken algebra ever..TY

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

right, thats what i got :)

jimthompson5910 (jim_thompson5910):

it should be \[\{x \neq \frac{8}{5} \}\]

jimthompson5910 (jim_thompson5910):

sry the symbol didn't show up, but you got the idea

OpenStudy (anonymous):

i wonder why it would have [ ] instead of { }?

OpenStudy (anonymous):

yes.

jimthompson5910 (jim_thompson5910):

typo perhaps

OpenStudy (anonymous):

haha, i get those at times

jimthompson5910 (jim_thompson5910):

lol that's why it's important to not always trust the computer

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

i got one more problem, its a word problem..i will post a new one, its been mind boggling for me

jimthompson5910 (jim_thompson5910):

k